OpenStudy (anonymous):

Subtract the complex expressions, and simplify. (-1.9 + 8.5i) – (-9.5 – 0.4i)

OpenStudy (anonymous):

\[\LARGE (a+b)-(-c-d)=a+b+c+d\] Remember: \[\LARGE (-)\cdot (-)=(+)\]

OpenStudy (anonymous):

@Notalgebrasmart do you get it ? ...

OpenStudy (anonymous):

7.6+8.9i

OpenStudy (anonymous):

= -1.9+9.5+8.5i+0.4i =7.6 + 8.9i

OpenStudy (anonymous):

all these formulas are kicking my tail krishnik so you just add them all up right.

OpenStudy (anonymous):

Yes. add real part separtately and comlex part separately

OpenStudy (anonymous):

makes alot of sense to do that. thanks shivam

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

lol, I apologize, It wasn't my intention to get you confused, I just used letters instead of numbers. :(. and yes they've given you the right answer ;).

OpenStudy (anonymous):

lol. i did get what u was saying. algebra just got so many formulas.